# Exploring Missing Modality in Multimodal Egocentric Datasets

### v1

@@ 1 Introduction
Multimodal video understanding has been the de facto approach for analyzing egocentric videos. Recent works have shown that the complimentary multisensory signals in egocentric videos are superior for understanding actions [32, 20, 21, 22, 28] and localizing moments [38, 2, 36]. However, multimodal systems need to be practical for real-world applications that could suffer from the incompleteness of modality inputs due to privacy, efficiency, or simply device failures [25]. For example, when predicting in real-time using a wearable device, parts of the recordings might be scrapped to preserve the privacy of the bystanders/camera wearer [14]. Furthermore, using all sensors could be expensive for a wearable device, opting for cheaper modalities such as audio or IMU. Thus, studying the impact of missing modalities is crucial for realistic performance expectations.
@@
Figure 1: Most commonly, we train the multimodal models on modal-complete data. These models fail when testing data has inputs with missing modalities (orange plots). Our proposed adaptation to missing modality significantly improves the performance across datasets (green plots). When all test inputs are modal-incomplete ($r_{test}=100\%$), we surpass unimodal performance (blue plot) by 5 points in Epic-Kitchens, and we are on par with the unimodal accuracy in Epic-Sounds.
@@
Still, the current effort to study the impact of missing modalities in egocentric datasets is limited. Most methods assume that all modality inputs are complete at training and inference time. Recent works have studied the effect of missing modalities for different tasks varying from recommendation systems to emotion recognition [33, 39, 31, 41, 24, 35]. Most works related to missing modalities have studied the problem at test time only [39, 35, 43, 41, 33, 3], while just a handful studied missing modalities in both training and testing time [24, 25, 30]. Similar to our setting, Lee et al. [25] propose a strategy to learn prompts for pre-trained backbones to deal with missing modalities. However, they analyze their method for image and text datasets only; we implement our version for action recognition and use it as a baseline in Section 4. More recently, Gong et al. [12] proposed a benchmark for multimodal generalization, focusing on few-shot learning recognition while considering missing modalities. Though the latter work proposes an interesting benchmark that includes a zero-shot and few-shot setup, no works have diagnosed how recent transformer-based approaches perform when modalities are missing for the action recognition setting.
@@
In this work, we study the problem of missing modalities in egocentric action recognition. First, we investigate how current transformer-based models are affected by incomplete modalities at test time. In Figure 1, we observe how the current state-of-the-art audiovisual recognition model, Multimodal Bottleneck Transformer (MBT) [32], trained on modal-complete inputs, suffers from a critical degradation in performance when the missing modality rate increases. The advantage of the multimodal backbone (orange) is lost when the missing modality rate in the test set exceeds $\sim 30\%$ (Epic-Sounds) and $\sim 70\%$ (Epic-Kitchens), at which point the unimodal model (blue) becomes a better alternative. To address this problem, we propose learning the missing modality ”template” during training to replace missing modalities at test time. We call this template the Missing Modality Token (MMT) and explain how to learn it in Section 3.4. Figure 1 (Epic-Kitchens) also shows how our approach (green) dramatically improves the test accuracy and stays at least 5% points above the unimodal performance even when the test-set is fully modal-incomplete.
@@
We verify the effectiveness of our method in 3 egocentric datasets, including Ego4D [14], which has a full coverage of RGB video but less than $70\%$ of the videos have audio. We extensively analyze our proposed training strategies, showing how to train with MMT under different missing modality scenarios. Our experiments show that our simple yet effective approach proposes a strong solution to this problem. Our contributions are threefold: (1) We present a thorough study of the missing modality problem in egocentric action recognition by considering effectiveness, efficiency, and the impact of the fusion strategy. (2) We propose Missing Modality Token (MMT) to deal with missing modalities at train and test time for egocentric video understanding. We propose two training strategies, namely fixed-replace and random-replace. (3) We extensively study our method and find that it significantly improves the presented baselines. Our work brings insights and paves the way to build multimodal backbones robust to missing modalities.
@@ 2 Related Work
Missing Modality Modeling. Addressing missing modalities presents a notable challenge, explored through various strategies by researchers from different areas. From medical applications [1] to sentiment analysis [3], missing modalities are a long-standing problem in multimodal understanding. Similarly, Ma et al. [30] and Colombo et al. [3] investigate missing modalities within a Bayesian Meta-learning framework. Meanwhile, Tsai et al. [39] and Zhao et al. [43] address unexpected missing or noisy modalities at test time using generative models to recreate missing inputs. Neverova et al. [33] focus on multimodal gesture recognition, employing depth, audio, and video streams, and updating network parameters based on different modality combinations. Our approach takes inspiration from [33] but uses a simpler learning approach that does not require different modality loss combinations. We learn a Missing Modality Token (MMT), which serves as a template for missing modalities during training without requiring complex pipelines using generative models.
@@
Other studies utilizing transformers [40], such as the work of Parthasarathy et al. [35], explore missing modalities at test time and propose training-time augmentations. Radevski et al. [37] apply multimodal distillation to enhance training for models using only RGB at test time. Ma et al. [31] develop strategies for optimal fusion layers and class tokens in the context of missing modalities, focusing on image-text datasets. Our research differs by demonstrating effectiveness across various fusion layers (Section 4.5), indicating that optimization of this parameter is not crucial. Lee et al. [25] proposes to learn to prompt large multimodal backbones for image and text classification when modalities are missing at train and test time. We adapt their method to our setting and show that ours is more practical and effective for dealing with missing modalities in egocentric videos. Lastly, Gong et al. [12] introduce a benchmark for handling missing modalities within the Ego4D dataset, tailored for few-shot classification  [^1] Code and data are not available. Though the previous approach proposes an interesting benchmark it introduces an extra layer of complexity by including few-shot and zero-shot recognition into the missing modality problem. Our work proposes to diagnose and study the problem in a simpler setting to understand the effect of missing modalities in standard egocentric video understanding.
@@
Multimodal Egocentric Video Understanding. Egocentric perception faces distinct challenges compared to traditional video understanding benchmarks such as ActivityNet [7] and Kinetics [19]. The nature of how egocentric datasets are captured means that they usually feature strongly aligned and synchronized audiovisual signals. Key benchmarks in this field, including Epic-Kitchens [4], Epic-Sounds [18], and the more recent and extensive Ego4D [14], have demonstrated the importance of audiovisual learning for understanding egocentric videos due to the complementary nature of the audio and visual modalities [38, 21, 20]. These datasets have facilitated the creation of several audiovisual backbones tailored for video understanding. Xiao et al. [42] introduced a CNN-based dual-stream architecture, utilizing SlowFast networks for the visual component [8] and a separate stream for audio [22]. With the advent and adaptability of transformer architectures, several studies have treated different modalities as input tokens for a multimodal transformer encoder [26, 23, 27, 10]. However, self-attention mechanisms can become prohibitively expensive as the number of tokens increases. To address this, Nagrani et al. [32] offered an efficient modality fusion technique that avoids costly self-attention in their Multimodal-Bottleneck Transformer (MBT). We build atop MBT and propose a method to make it robust for missing modalities at train and test time.
@@ 3.1 Problem Statement, Setup, and Evaluation
Given the training and testing multimodal data samples, let us denote the missing modality rates in each set with $r_{train}$ and $r_{test}$, respectively. These rates are computed by dividing the number of modal-incomplete samples by the total number of samples. Note that we only consider the setting of one modality being missing in the dataset. Our work discusses the strategies of training a multimodal model under two scenarios: training-with-missing, i.e. some samples in the training set are modal-incomplete ($r_{train}\neq 0\%$), or training-with-complete i.e. all samples in the training set all modal-complete ($r_{train}=0\%$).
To observe the trained model’s behavior under different missing modality severity levels, we create several variants of the test set by manually removing the modality information from the samples until $r_{test}=100\%$. For non-adapted models, we replace missing modality inputs at test time with dummy (empty) tensors.
@@ 3.2 Efficient and Effective Multimodal Fusion
We deal with missing modalities while considering the methods proven to be the most effective for multimodal fusion. Previous transformer-based methods addressing missing modalities looked mostly at basic methods, such as early or mid-fusion with cross-modal self-attention, where all tokens are concatenated at the fusion layer. This does not scale well in videos due to the attention mechanism’s quadratic complexity (to the input size) [32]. Instead, we use the current state-of-the-art audiovisual fusion, MBT [32], which proved to be more efficient and effective. The bottleneck transformer in the MBT design allows the model to distill and propagate the most essential information across modalities where each modality performs self-attention only with a small number of learnable ”bottleneck” tokens. Such design is especially useful for information-dense (redundant) modalities like video.
Architecture Input dummy tokens Audio $r_{test}$ Accuracy
Self-attention
0% 45.3%
50% 32.7%
75% 26.7%
100% 20.5%
Bottleneck
0% 55.5%
50% 34.3%
75% 23.7%
100% 13.6%
Bottleneck
0% 55.5%
50% 38.7%
75% 30.2%
100% 21.7%
Video-only
0% - 100% 41.4%
Table 1: Bottleneck fusion vs. vanilla self-attention with early fusion in Epic-Sounds dataset. Bottleneck fusion is shown with 2 ways of manipulating missing inputs at test time: passing dummy tokens or removing the tokens. All shown models fall far behind the unimodal (video) performance.
@@
Figure 2: Learning with MMT explained. Left: Given modal-incomplete data, it is still unclear how to effectively train a multimodal model and then predict with it. Right: We demonstrate how we perform training and testing with modal-incomplete data by using Missing Modality Token (MMT). MMT learns the best representation of missing inputs. At training time, we use both modal-incomplete and modal-complete samples to train MMT. With the latter, we randomly decide for which samples to replace the signal tokens with MMT. This strategy allows us to effectively represent the missing modality tokens at test time.
@@
Fusion Layer. A key aspect of a multimodal fusion approach is the design of the fusion layer ($L_{f}$). $L_{f}$ is the layer at which the cross-modal interactions happen. We observe the performance of the original bottleneck model with modal-complete audiovisual inputs and train the model with different fusion layers. We show the test accuracy for these models in Figure 6 (marked as Non-Adapted) for the Epic-Kitchens and Epic-Sounds datasets. We find that the performance does not change significantly with different fusion layers when $r_{test}=0$. However, the fusion layer does make a difference when inputs are incomplete (e.g., 35% test accuracy with $L_{f}=0$ vs. 42% with $L_{f}=11$ in Epic-Sounds at $r_{test}=50\%$), as shown in Section 4.5. Overall, fusing earlier is preferred in Epic-Kitchens, but fusing later gives better results in Epic-Sounds. This outcome is consistent with the observations from the previous work [31]: the best fusion strategy is dataset-specific. While this observation might be intuitive, it is impractical when the model is very sensitive to the fusion layer, as searching for the best layer might be computationally expensive. Thus, we analyze the effect of the fusion layer when modalities are missing and show the effect of our approach.
3.3 Model’s intrinsic capabilities with missing modalities
Before jumping to any training adaptation strategy, we first explore the capability of the models trained without any specific adaptation to the missing modality.
Bottlenecks Condensing the Information. As mentioned above, MBT uses a small set of learnable fusion tokens to exchange information between the modalities. One might wonder if the ability of the bottleneck to ”condense” the information already plays some role in dealing with missing inputs. To investigate this, Table 1 compares training a vanilla self-attention model with $L_{f}=0$ (row 1) with the training MBT with the same fusion layer (row 2). Table 1 shows that vanilla self-attention performs poorly in the multimodal setup with complete modalities, losing 10 points, compared to the bottleneck fusion. Surprisingly, when the modality inputs are incomplete, vanilla self-attention performs better (20% vs 13.6% with $r_{test}=100\%$). This shows that while the bottlenecks are effective for multimodal fusion, they are sensitive to missing modalities and they must be adapted to address incomplete modalities.
Only Passing Complete Inputs at Test-Time. As an alternative to passing an empty tensor when the modality is missing, one could suggest simply discarding the tokens from the missing signal and only feeding the transformer with the non-missing modality. The previous alternative is only possible when using transformers, as they can take a variable number of input tokens. Table 1 shows that this strategy (row 3) performs better than just passing empty tensors (row 2). Nevertheless, the model fails to reach video-only performance (row 4), even with a lower missing rate $r_{test}=50\%$, and only reaches 21.7% when $r_{test}=100\%$. Thus, this adaptation is still far from optimal, and we seek better solutions.
@@ 3.4 Our Approach to Dealing with Missing Modalities
We suggest a simple and generic way to deal with missing modalities. Instead of passing dummy inputs or discarding the tokens of the missing inputs at test time, we propose to learn a ”template” for the missing inputs. We introduce a learnable Missing Modality Token (MMT), which is trained to represent the tokens of the missing modality by learning from the tokens of the non-missing modality. As shown in Figure 2 (left) and explained in the previous sections, it is unclear how to deal with inputs with missing modalities. To this end, we introduce the Missing Modality Token (MMT). Figure 2 (right) illustrates how the MMT replaces the token of missing modalities. Our solution is parameter efficient since MMT is shared across samples. However, we still add positional embeddings to each token to preserve the location information of each missing patch. As shown in Figure 2 the MMT is learned at training time by alternating complete-input tokens with MMT. We propose two strategies to train with MMT: fixed-replace and random-replace and discuss how to apply them in training-with-complete and training-with-missing scenarios. At test time, we replace the missing modality inputs with our learned MMT.
Strategy fixed-replace. This strategy assumes that there is a fixed subset of the training instances with a missing modality. MMT tries to learn the best representation of the missing modality using this fixed subset. This strategy can be straightforwardly applied for training-with-missing scenario, as $r_{train}$ of the training samples are modal-incomplete. In training-with-complete, we manually create this fixed subset of modal-incomplete instances by scraping the modality data in some instances for the whole training process. Essentially, we convert this scenario to training-with-missing by making $r_{train}>0$. We input MMTs to the transformer to represent the tokens of the scrapped inputs. Note that when training-with-missing we can still control $r_{train}$ by removing the modality from more instances atop the ones that are naturally missing. In this strategy, we completely discard these samples’ modality information during the whole training.
Strategy random-replace. Instead of dropping the modality in a fixed subset of training samples, we remove the modality for the random samples (keeping $r_{train}$ unchanged). For each training sample, with probability $p$, the ”missing” modality tokens will be replaced with MMT. Our intuition is that by picking the samples randomly, rather than using a fixed pool, the model can ”see” the same samples in both missing and complete modality scenarios. By doing so, the model can learn a better representation of the missing inputs. In training-with-missing, there is still a fixed set of modal-incomplete training samples. We input MMTs for these samples, too. Basically, in this scenario, fixed-replace is given, and we randomly drop from the modal-complete subset.
@@
Inference. Regardless of the strategy used, we replace the missing inputs tokens with the learned MMT at test time.

### v2

@@ Abstract
Keywords:
Missing Modality Multimodal Video Recognition Egocentric Videos
@@ 1 Introduction
Multimodal video understanding has been the de facto approach for analyzing egocentric videos. Recent works have shown that the complimentary multisensory signals in egocentric videos are superior for understanding actions [37, 25, 24, 26, 33] and localizing moments [43, 2, 41, 47]. However, multimodal systems need to be practical for real-world applications that could suffer from the incompleteness of modality inputs due to privacy, efficiency, or simply device failures [30]. For example, when predicting in real-time using a wearable device, parts of the recordings might be scrapped to preserve the privacy of the bystanders/camera wearer [16, 14]. Furthermore, using all sensors could be expensive for a wearable device, opting for cheaper modalities such as audio or IMU [17]. Thus, studying the impact of missing modalities is crucial for realistic performance expectations.
@@
Figure 1: Most commonly, we train the multimodal models on modal-complete data. These models (orange) fail when encountering modal-incomplete data at test time. Our proposed adaptation to the missing modality (green) significantly improves the performance across datasets. When all test inputs are modal-incomplete ($r_{test}=100\%$), we surpass unimodal performance (purple) by 5 points in Epic-Kitchens, and double the baseline performance in Ego4D-AR.
@@
Still, the current effort to study the impact of missing modalities in egocentric datasets remains rather limited. Most methods presume all modal inputs to be intact during training and inference. Recent works have studied the effect of missing modalities for different tasks varying from recommendation systems to emotion recognition [38, 44, 35, 46, 29, 40]. Notably, the majority of research concerning missing modalities has primarily addressed the issue during testing [44, 40, 50, 46, 38, 3], while just a handful studied it across both training and testing phases [29, 30, 36]. Similar to our setting, Lee et al. [30] propose a strategy to learn prompts for pre-trained backbones to deal with missing modalities. However, they analyze their method for image and text datasets only; we implement our version for action recognition and use it as a baseline in Sec. 4. More recently, Gong et al. [14] proposed a benchmark for multimodal generalization, focusing on few-shot learning recognition while considering missing modalities. Though the latter work proposes an interesting benchmark that includes a zero-shot and few-shot setup, no works have diagnosed how recent transformer-based approaches perform when modalities are missing for the action recognition setting.
@@
In this work, we study the problem of missing modalities in egocentric action recognition. First, we investigate how current transformer-based models are affected by incomplete modalities at test time. In Fig. 1, we observe how the current state-of-the-art audiovisual recognition model, Multimodal Bottleneck Transformer (MBT) [37], trained on modal-complete inputs, suffers from a critical degradation in performance when the missing modality rate increases. The advantage of the multimodal backbone (orange) is lost when the missing modality rate in the test set exceeds $\sim 27\%$ (Ego4D-AR) and $\sim 70\%$ (Epic-Kitchens). At this point, the unimodal model (purple) becomes a better alternative. To address this problem, we propose learning the missing modality "template" during training to replace missing modalities at test time. We call this template the Missing Modality Token (MMT) and explain how to learn it in Sec. 3.4. Fig. 1 (Epic-Kitchens) also shows how our approach (green) dramatically improves the test accuracy and stays at least 5 points above the unimodal performance even when the test set is fully modal-incomplete. Furthermore, our methods enable better multimodal performance overall in modal-incomplete Ego4D-AR.
@@
We verify the effectiveness of our method in 3 egocentric datasets, including Ego4D [16], which has a full coverage of RGB video, but only $70\%$ of videos have audio. We extensively analyze our proposed training strategies, showing how to train with MMT under different missing modality scenarios. Our experiments show that our simple yet effective approach proposes a strong solution to this problem. Our contributions are threefold: (1) We present a thorough study of the challenge of missing modalities in egocentric action recognition. This involves exploring datasets with varying degrees of modal incompleteness and assessing the influence of the fusion layer. (2) We propose the Missing Modality Token (MMT) as a novel solution to address missing modalities during both the training and testing phases. Additionally, we propose a training strategy, termed random-replace, to enhance the efficacy of models utilizing MMT. (3) We extensively evaluate our method and demonstrate its notable improvement over existing baselines. Through our work, we provide valuable insights and lay the groundwork for developing multimodal backbones that exhibit robustness in the face of missing modalities.
@@ 2 Related work
Addressing missing modality. Addressing missing modalities presents a notable challenge, explored through various strategies by researchers from different areas. From medical applications [1] to sentiment analysis [3], missing modalities are a long-standing problem in multimodal understanding. Some methods [42, 12, 11] distill the knowledge from a multimodal teacher to an unimodal RGB model. Others are tailored for scenarios where test data is multimodal yet incomplete in terms of modalities. For example, Ma et al. [36] and Colombo et al. [3] investigate missing modalities within a Bayesian Meta-learning framework. Meanwhile, Tsai et al. [44], Zhao et al. [50], and Woo et al. [48] attempt to reconstruct missing inputs. Neverova et al. [38] focus on multimodal gesture recognition, employing depth, audio, and video streams, and updating network parameters based on different modality combinations. Most of these works rely on modality-specific architectures [28, 20] and/or use complex generative pipelines. Our approach uses a generic multimodal Transformer [45]. Furthermore, these methods assume that the training data is fully modal-complete, which is not the case in current large-scale datasets [17]. Instead, our method applies to modal-complete and modal-incomplete training sets.
@@
Recent studies utilizing transformers, such as the work of Parthasarathy et al. [40], explore missing modalities at test time and propose training-time augmentations. Ma et al. [35] develop strategies for optimal fusion layers and class tokens in the context of missing modalities, focusing on image-text datasets. Our research differs by demonstrating the effectiveness of our method across various fusion layers (Sec. 4.7), which avoids any expensive fusion policy learning. Lee et al. [30] proposes to learn to prompt large multimodal backbones for image and text classification when modalities are missing at train and test time. We adapt their method to our setting and show that ours is more practical and effective for dealing with missing modalities in egocentric videos. Lastly, Gong et al. [14] introduce a benchmark for handling missing modalities within the Ego4D dataset, tailored for few-shot classification  [^1] Code and data are not available. Our work proposes to diagnose and study the problem in a simpler setting to understand the effect of missing modalities in egocentric video understanding. We want to note that most of related transformer-based methods [14, 35, 40] do not provide the code, which makes it challenging to compare to.
@@
Multimodal egocentric video understanding. Egocentric perception faces distinct challenges compared to traditional video understanding benchmarks such as ActivityNet [7] and Kinetics [23]. The nature of how egocentric datasets are captured means that they usually feature strongly aligned and synchronized audiovisual signals. Key benchmarks in this field, including Epic-Kitchens [4], Epic-Sounds [22], and the more recent and extensive Ego4D [16], have demonstrated the importance of audiovisual learning for understanding egocentric videos due to the complementary nature of the audio and visual modalities [43, 24, 25]. These datasets have facilitated the creation of several audiovisual backbones tailored for video understanding. Xiao et al. [49] introduced a CNN-based dual-stream architecture, utilizing SlowFast networks for the visual component [8] and a separate stream for audio [26]. With the advent and adaptability of transformer architectures, several studies have treated different modalities as input tokens for a multimodal transformer encoder [31, 27, 32, 10]. However, self-attention mechanisms can become prohibitively expensive as the number of tokens increases. To address this, Nagrani et al. [37] offered an efficient Multimodal Bottleneck Transformer (MBT) that avoids costly self-attention. We build atop MBT and introduce Missing Modality Token to make it robust for missing modalities at train and test times.
@@ 3.1 Problem statement, setup, and evaluation
Given the training and testing multimodal data samples, let us denote the missing modality rates in each set with $r_{train}$ and $r_{test}$, respectively. These rates are computed by dividing the number of modal-incomplete samples by the total number of samples. Note that we only consider the setting of one modality being missing in the dataset. Our work discusses the strategies of training a multimodal model under two scenarios: modal-incomplete training set i.e., some samples in the training set are modal-incomplete ($r_{train}\neq 0\%$), or modal-complete training set i.e., all samples in training set all modal-complete ($r_{train}=0\%$).
To observe the trained model’s behavior under different missing modality severity levels, we create several variants of the test set by manually removing the modality information from the samples until $r_{test}=100\%$.
@@ 3.2 Efficient and effective multimodal fusion
We deal with missing modalities while considering the methods proven to be the most effective for multimodal fusion. Previous transformer-based methods addressing missing modalities looked mostly at basic methods, such as early or mid-fusion with cross-modal self-attention, where all tokens are concatenated at the fusion layer. This does not scale well in videos due to the attention mechanism’s quadratic complexity (to the input size) [37, 30]. Instead, we use the current state-of-the-art audiovisual fusion, MBT [37], which proved to be more efficient and effective. The bottleneck transformer in the MBT design allows the model to distill and propagate the most essential information across modalities where each modality performs self-attention only with a small number of learnable "bottleneck" tokens. Such design is especially useful for information-dense (redundant) modalities like video.
@@
Figure 2: Learning and Predicting with Missing Modalities. Left: Given modal-incomplete data, it is still unclear how to effectively train and predict with a multimodal model (we present some naive baseline methods in Sec. 3.3). Right: To address this issue, we introduce a Missing Modality Token (MMT). During training, MMT learns the representation of missing inputs from modal-incomplete samples and modal-complete samples. For the latter, we use random-replace to let the network observe the missing inputs and thus learn better representations (Sec. 3.4). At test time, we replace the tokens of missing inputs with MMT to effectively represent them.
@@
Fusion Layer. A key aspect of a multimodal fusion approach is the design of the fusion layer $L_{f}$. $L_{f}$ is the layer at which the cross-modal interactions happen. We observe the performance of the original bottleneck model with modal-complete audiovisual inputs and train the model with different fusion layers. We show the test accuracy for these models in Figure 6 (marked as baseline) for the Epic-Kitchens and Epic-Sounds datasets. We find that the performance does not change significantly with different fusion layers when $r_{test}=0$. However, the fusion layer does make a difference when inputs are incomplete (e.g., 35% test accuracy with $L_{f}=0$ vs. 42% with $L_{f}=11$ in Epic-Sounds at $r_{test}=50\%$), as shown in Section 4.7. Overall, fusing earlier is preferred in Epic-Kitchens, but fusing later gives better results in Epic-Sounds. This outcome is consistent with the observations from the previous work [35]: the best fusion strategy is dataset-specific. While this observation might be intuitive, it is impractical when the model is very sensitive to the fusion layer, as searching for the best layer might be computationally expensive. Thus, we analyze the effect of the fusion layer when modalities are missing and show the effect of our approach in Sec. 4.7.
3.3 Intuitive baselines
While designing effective solutions for missing modalities is challenging (Fig. 2), one could suggest simple and intuitive ways of adapting to missing inputs at test time. We propose the following training-free baselines to deal with missing modality at inference time.
Passing missing inputs as tensors with zeros. We employ a straightforward approach of substituting missing modality inputs with tensors filled with zeroes. This method, widely acknowledged in the literature [1, 30, 40], is favored for its simplicity and ease of implementation in practice. Thus, unless stated otherwise, we adopt this method as our primary baseline and refer to it as baseline.
Only pass complete inputs. As transformers can process sequences of varying lengths, we can selectively omit tokens corresponding to missing signals and exclusively supply the transformer with non-missing modality tokens. While intuitive, this approach becomes less practical when inference involves batch sizes greater than 1 and not all inputs within the batch exhibit modal incompleteness.
@@ 3.4 Our approach to dealing with missing modalities
We suggest a simple and generic way to deal with missing modalities. Instead of passing tensors filled with zeroes or discarding the tokens of the missing inputs at test time, we propose to learn a "template" for the missing inputs. We introduce a learnable Missing Modality Token (MMT), which is trained to represent the tokens of the missing modality by learning from the tokens of the non-missing modality.
Figure 2 (right) illustrates how MMT represents the tokens of the missing modality at train and test time. When the model encounters a modal-incomplete sample, MMT, a shared parameter, is repeated to match the missing modality token number and passed to the transformer, along with the tokens of non-missing modalities. Note that each token has a different positional embedding added, similar to the mask token in [18]. With this intuitive way, MMT observes all modal-incomplete samples and leverages the non-missing modality tokens to learn.
If trained with modal-incomplete samples only, the model never encounters the inputs it tries to mimic. Furthermore, MMT is limited in the number of training samples (e.g. in a dataset with $r_{train}=10\%$, MMT will only encounter $10\%$ of the data). We try to facilitate it and suggest random-replace strategy. With random-replace, input tokens of modal-complete samples can also be used to train MMT. Namely, for each modal-complete training sample, with probability $p$, the tokens of one modality will be replaced with MMT. When $p=0$, the MMT learns with naturally missing inputs only. If $p$ is set to a non-zero value, the strategy provides more training samples for learning MMT and lets the network observe the same sample as modal-incomplete and modal-complete; thus encouraging the model to understand the relationships and dependencies of the modalities. However, setting $p$ too high might cause high information loss and hinder the performance, especially if we drop the tokens of a more information-dense modality, as we will show in Sec. 4.4
Thus, there are 2 ways of providing training samples for MMT: (1) Using samples with naturally missing modalities and (2) Randomly replacing the tokens of complete samples with MMT. Below, we discuss how both ways are used with modal-complete and modal-incomplete training sets.
Modal-incomplete training set. Recall that $r_{train}$ of the training samples are modal-incomplete and $100\%-r_{train}$ are modal-complete. Thus, MMT can:
1. Learn from modal-incomplete only. During training, we use MMT to represent the missing modality inputs for modal-incomplete samples. The tokens of modal-complete samples are never replaced with MMT i.e., $p=0$.
2. Learn from modal-complete and modal-incomplete. Use modal-incomplete samples as in (1), and random-replace with non-zero $p$.
Modal-complete training set. As all training samples have complete multimodal inputs, MMT is trained using random-replace with $p>0$.
@@
Inference. Regardless of the strategy, we replace the missing input tokens with the learned MMT at test time.